OpenStudy (anonymous):

PLEASE HELP ME ON THIS ALGEBRA PROBLEM:-)Rewrite in simplest rational exponent form square root x multiplied by 4 square root x . Show each step of your process.

OpenStudy (anonymous):

@Vocaloid @veechan @Gambino @goformit100 @Destinyyyy @dan815

OpenStudy (anonymous):

@LiveLaughDie @LiteNing1337 @LexiLuvv2431

OpenStudy (anonymous):

Can you use the equation button to correctly type out the equation so its easy to read :) I can help you, I just want to make sure that my information is accurate.

OpenStudy (anonymous):

okay! thanks

OpenStudy (anonymous):

\[\sqrt{x} \times \sqrt[4]{x}\]

OpenStudy (anonymous):

and then in simplest rational exponent form :-) @LexiLuvv2431

OpenStudy (anonymous):

Okayy, do you have any answer choices? or is this a word problem

OpenStudy (anonymous):

its just a question where I need to explain each step and how i got it and the answer of course :)

OpenStudy (anonymous):

Okayy so, I'm gonna walk you through this just so you know for next time :) Do you know where to begin?

OpenStudy (anonymous):

not really for this one

OpenStudy (anonymous):

You're going to rewrite \[\sqrt{x}\] with rational exponents first. so \[\sqrt{x}\] = \[x_{2}^{1}\]

OpenStudy (anonymous):

where'd you get the 2 from?

OpenStudy (anonymous):

That is x with an exponent of 1/2 if you're confused.

OpenStudy (anonymous):

okay thanks!

OpenStudy (anonymous):

2 is a rational exponent of x thats were the 2 comes from

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

Next we have \[\sqrt[4]{x}\] You then take the rational exponent from here too which is the 4 aand turn it into a fration so it would look like this \[x_{4}^{1}\] x has the exponent of 1/4

OpenStudy (anonymous):

that makes sense

OpenStudy (anonymous):

So the final product is